A convergent (biconvex) lens is placed inside a jar filled with a liquid. The lens has focal length 20cm, when in air and its material has a refractive index of 1.5 . If the liquid has a refractive index of 1.6, the focal length of the lens while in the jar, is:-

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Student
 
When the lens is in air, we have (from lens maker’s equation),
1/20 = [(1.5/1) – 1] (1/R1- 1/R2).
We are not bothered about the signs of R1 and R2 since they are unknown quantities.
Even though we know that the convergent lens will become divergent in the denser medium, we write lens maker’s equation without bothering about the sign of its unknown focal length in the liquid. If ‘f’ is the focal length in the liquid, we can write,
1/f = [(1.5/1.6) – 1] (1/R1 – 1/R2).
Dividing the first equation by the second, we obtain f/20 = 0.5×1.6/(-0.1) from which f = -160 cm 
 
Regards
Arun (askIITians forum expert)